Suppose you invest $7000 at an annual rate of 6.25% compounded continuously. How much time will it take for the investment to reach a value of $56000?
Answer by stanbon(75887)

Suppose you invest $7000 at an annual rate of 6.25% compounded continuously. How much time will it take for the investment to reach a value of $56000
--------------------------
A(t) = P*e^(rt)
56000 = 7000*e^(0.0625t)
e^(0.0625t) = 8
Take the natural log to get:
0.0625t = ln(8)
t = 33.27 years
=========================
Cheers,
Stan H.